I am thinking of three consecutive negative numbers. If I multiply the first with the second and then subtract three times the third, the result is 57. What are the numbers?
Answer by sowmya(32)

I am thinking of three consecutive negative numbers. If I multiply the first with the second and then subtract three times the third, the result is 57. What are the numbers?
solution:
Let the three numbers be x, x+1 and x+2
x(x+1)-3(x+2) = 57
x^2+x-3x-6=57
x^2-2x-63=0
x^2-9x+7x-63=0
(x+7)(x-9)=0
Hence x = -7 or x=9
But the three numbers shd be negative.So x= -7
Hence the three consecutive numbers are -7,-8,-9